OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ @Luigi0210 @AravindG

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

no thanx i need help not demands so... bye

OpenStudy (nurali):

f(n) = f(1) + f(2) + f(n - 1) f(3)=2+3+3=11 f(4)=2+3+11=16 f(5) = 2+3+16=21

OpenStudy (anonymous):

tnx

OpenStudy (nurali):

Anytime.